the random sample shown below was selected from a normal distribution
7, 6 , 5 , 6 , 9, 3
a.) construct a 90% confidence interval for the population mean u
(__,__)
Answer #1

From sample data

\bar{x} = \sum X / n = 6

S = sqrt [ \sum X2 - n \bar{x} 2 / n-1 ] = 2

t critical value at 0.10 significance with 5 df = 2.015

90% confidence interval for \mu is

\bar{x} - t * S / sqrt(n) < \mu < \bar{x} + t * S / sqrt(n)

6 - 2.015 * 2 / sqrt(6) < \mu < 6 + 2.015 * 2 / sqrt(6)

4.35 < \mu < 7.65

90% CI Is ( 4.35 , 7.65 )